# Security Weaknesses of Copilot Generated Code in GitHub

### v2

@@ 2.3. Security Static Analysis
Vulnerabilities detection is critical to improve software security and ensure quality  (Iannone et al., 2022). Vulnerability can be checked either statically or dynamically. Dynamic analysis techniques are more sound and precise but lack coverage (Ernst, 2003). On the other hand, static analysis is less precise but offers greater coverage and allows users to analyze programs without the need to execute them (Sui et al., 2020). OWASP (OWASP, [n. d.]) provides a list of commonly used static analysis tools for security analysis. This includes tools like CodeQL, a general-purpose automatic scanning tool; FindBugs, a tool for Java programs; ESLint, a tool for JavaScript programs; Bandit, a tool for Python programs; and GoSec, a tool for Go programs. Such tools have been widely used in previous security analysis research  (Tómasdóttir et al., 2018; Pearce et al., 2022; Siddiq et al., 2022).
Kaur et al. (Kaur and Nayyar, 2020) compared static analysis tools for vulnerability detection in scanning C/C++ and Java source code. Tomasdottir et al. (Tómasdóttir et al., 2018) conducted an empirical study on ESLint, the most commonly used JavaScript static analysis tool among developers. Pearce et al. (Pearce et al., 2022) used CodeQL to scan security weaknesses in the generated Python and C++ code. Siddiq et al. (Siddiq and Santos, 2022) used Bandit to check Python code generated using a test dataset. Lisa et al. (Do et al., 2020) reported on users’ goals, motivations, and strategies when using static analysis tools.
These static analysis tools support different analysis algorithms and techniques. By using multiple tools for analysis, potential weaknesses in the code can be discovered from different perspectives and levels, avoiding omissions and improving the accuracy of the analysis. Our study first used CodeQL to scan the collected code snippets. CodeQL is an open-source tool that supports multiple languages, including Java, JavaScript, C++, C#, and Python. It can find weaknesses in a codebase based on known weaknesses/rules. In addition, to obtain more comprehensive scan results, we supplemented the scan of code in different languages with static analysis tools (i.e., Cppcheck and Bandi) tailored to specific languages.
@@ 3. Research Design
In this section, we describe our research design in detail. In Section 3.1, we first define our Research Questions (RQs), followed by the process of collecting and filtering the code snippets generated by Copilot in Section 3.2. We then explain the security analysis performed on the identified snippets and the process of filtering the raw results generated by static analysis tools in Section 3.3.
@@ 3.1. Research Goal and Questions
This study aims to understand the potential security weaknesses in Copilot-generated code produced in real-world GitHub projects. We first collect code snippets generated by Copilot from GitHub projects as our data source, which is the largest source of open-source projects. It should be noted that it is not possible to access all the code generated by Copilot in GitHub projects, as there is no direct way to identify if part of a file was generated by Copilot (i.e., source files do not contain any signatures to indicate if Copilot generates the code). However, we can identify many code snippets by searching the repository description and the comments provided in the code (see the details in Section 3.2.2).
We chose to focus on Python and JavaScript code snippets as they are currently the two most popular programming languages used by developers according to a recent survey in 2022 by GitHub (oct, [n. d.]), and also the ones most frequently used with GitHub Copilot (Zhang et al., 2023). We first analyzed the functionality and application domains of the collected code snippets to get the demographic information of the Copilot-generated code. Next, we performed a security analysis on the identified code. We selected static analysis because of its coverage and the ability to analyze programs without executing them. Dynamic analysis is more sound and precise (as it can reason about the program behaviour) but also lacks coverage (Sui et al., 2020). Using static analysis will allow us to run the analysis on a segment of the program (smaller snippet) without needing the full program to be executed. Static security analysis tools have been widely used by developers and companies (Majdinasab et al., 2024; Ernst, 2003; Pearce et al., 2022). After obtaining the scan results, we manually checked the results to remove false positives reported by the static analysis tools. We finally used CWEs to classify the filtered results for further analysis to answer the RQs.
We conducted this empirical study by following the guidelines of Easterbrook et al. (Easterbrook et al., 2008). The RQs, their rationale, and the research process of this study (see Figure 1) are detailed in the subsections below.
@@
Rationale: Copilot may produce code suggestions that developers accept, but these suggestions may include security weaknesses that could make the program vulnerable. The answer to RQ1 helps understand the frequency of security weaknesses developers encounter when using Copilot in production.
RQ2. What security weaknesses are present in the code snippets generated by Copilot?
Rationale: Copilot-generated code may contain a variety of security weaknesses (Pearce et al., 2022), and developers should conduct a rigorous security review before accepting the generated code suggestions. Copilot’s documentation notes the following: “users of Copilot are responsible for ensuring the security and quality of their code (GitHub, 2023b)”. By identifying common or recurring weaknesses, developers can be more prepared to prevent, mitigate, or fix these security weaknesses. The answer to RQ2 can help developers better understand possible security weaknesses in the code generated by Copilot, allowing them to be vigilant about the generated code before it is integrated into their programs.
RQ3. How many security weaknesses belong to the MITRE CWE Top-25 security weaknesses?
Rationale: The MITRE list contains the Top 25 most dangerous security weaknesses (The MITRE Corporation, 2023a), offering a benchmark for gauging the severity of weaknesses in the Copilot-generated code (Pearce et al., 2022). The answer to RQ3 can help to understand whether the code generated by Copilot contains widely recognized types of security weakness and Copilot’s ability to handle these recent and common weaknesses.
@@ 3.2. Data Collection and Filtering
We chose GitHub as the primary data source for answering our RQs. GitHub is widely used by developers. As the world’s largest code hosting platform, GitHub contains millions of public code repositories and offers access to a large number of code resources, allowing us to cover multiple programming languages and project types in our study (Cosentino et al., 2017). We identify code snippets generated by Copilot from GitHub projects.
@@ 3.2.1. Code Snippets Collection
Step 1. To identify the code snippets generated by Copilot, we first conducted a pilot search to formulate our search keywords. First, we used “GitHub Copilot” and “Copilot” as our search keywords. As expected, we found that the term “Copilot” is a vague term, referring not only to the code generation tool launched by GitHub, but also to tools in the aviation or telemetry fields. Therefore, using the keyword solely may return irrelevant content that may not be related to the use of the tool. On the other hand, using “GitHub Copilot” can exclude content unrelated to the code generation tool Copilot and narrow the search scope, which is what we have used to locate the code snippets.
However, even with this basic search keyword, we still need to carefully filter the search results to ensure that they are truly related to GitHub Copilot. Although using “GitHub Copilot” increases the relevance of the results to Copilot, these results are not necessarily the code snippets generated by Copilot. It should be noted that many code snippets containing the “GitHub Copilot” keyword in the search results display GitHub Copilot as text. Developers may use them to describe their experience using Copilot to generate code or showcase information related to Copilot. These code snippets are not what we seek as they do not directly relate to the code generated by Copilot. Our target is code generated by Copilot, not code snippets containing the keyword “Copilot”.
Our observations from the pilot search showed that using keywords such as “by GitHub Copilot”, “use GitHub Copilot”, and “with GitHub Copilot” can improve the accuracy of search results. These keywords enable us to focus more on the code generated using Copilot rather than code snippets that contain other content related to Copilot. In addition, since Python and JavaScript are the two most popular programming languages used by developers, and also the most frequently used with GitHub Copilot (see Section 3.1), we further limited the types of code snippets during the search to Python and JavaScript. We collected the search results under the Code label. Considering that some projects declare using GitHub Copilot generated code in their README files or project description provided in GitHub, we decided to retain the results from the Repository label in the search results. Figure 2 shows an example of our search process.
@@
Table 1 reports the search terms we used and the number of search results obtained from GitHub. In this step, we collected a total of 8,157 results, of which 7,749 were from the Code label, and 408 were from the Repository label. The same search result may contain multiple keywords, which means that there are duplicate projects in the collected data. After removing duplicate projects, we obtained a total of 2023 search results, of which 1847 were from the Code label, and 176 were from the Repository label.
Figure 2. Example of the search process
Table 1. Search results based on different terms used
# Search Term # Code # Repositories
1 “By GitHub Copilot” 2549 94
2 “Use GitHub Copilot” 1822 123
3 “With GitHub Copilot” 3378 191
Total 7749 408
@@ 3.2.2. Filtering Code Snippets
Step 2. After obtaining the data from the keyword searches, we further filtered them by not only considering the accuracy of the keywords but also by investigating the project’s documentation, code comments, and other metadata in the search results to determine whether they were generated by GitHub Copilot. Additionally, since we aimed to obtain code snippets used in real-world projects, we excluded search results used to solve simple programming practice problems on platforms, such as bisection lookup and fast sorting tasks from LeetCode. We consider these programming exercises to be more closely related to correctness than security.
@@
We first conducted a pilot data filtration to better filter the search results. We begin by explaining the terminology used in data filtering: the search results under the Repository label are the projects that contain code files, and the search results under the Code label are individual code files. Those code files contain code snippets generated by Copilot. In filtering the projects, we followed three criteria including two inclusion criteria and one exclusion criterion. Inclusion Criterion 1: for search results under the Repository label, we identified projects that are fully generated by Copilot, as declared in the project description or the associated README file(s). We retained code files for Python and JavaScript. Inclusion Criterion 2: For search results under the Code label, we retained code files with comments showing the code generated by Copilot. Exclusion Criterion 1: We excluded code files used to solve simple programming practice problems. We provide examples for the three criteria in Figure 3, Figure 4, and Figure 5. As shown in the example in Figure 3 for the Repository label, we kept all the Python files. In the next example in Figure 4, we kept the entire file where the Copilot-generated code snippet was located. In Figure 5, the code snippet was removed as it was determined that the code just solved a simple algorithmic problem.
Specifically, the pilot data filtering process consisted of the following steps: (1) The first author randomly selected 40 projects and 400 code files from the search results from the Repository label and the Code label, respectively. (2) Two authors independently labelled whether Copilot generated these projects and code files. (3) The first author compared the labelling results by the two coders and calculated the level of agreement between them using the Cohen’s Kappa coefficient (Cohen, 1960). (4) If there were any results they were unsure of or disagreed with, the two authors discussed until they reached an agreement. The Cohen’s Kappa coefficients were 0.79 for the projects (from the Repository label) and 0.85 for the code files (from the Code label), which were both higher than 0.7, indicating a high level of agreement between the two coders and ensuring a good accuracy of the labelling results. After completing the pilot data labelling, the first author checked the rest of the search results (1,447 code files and 136 projects). After removing the duplicate results, we retained 79 projects under the Repository label and 180 code files under the Code label. Step 3. To further get the application domains represented in the dataset, we categorized the projects based on the project’s description documents and the specific function of the code. Projects were classified into one of the following categories: Games, Web Applications, Utility Tools, AI Applications, Network Communication Applications, and Others. Figure 6 shows the classification and distribution of the application domains of the Copilot-generated code in the dataset from the Repository and Code labels respectively.
For the code files retained under the Repository label, we consider the entire code file as code generated by Copilot. In other words, we assume that Copilot generates all the code in the file because it was stated in the README file that it was all generated by Copilot. For code files retained under the Code label, we know that the files contain code snippets, perhaps even just a few lines of code, generated by Copilot. Instead of identifying the specific Copilot-generated code in this step, we combine the warning messages from the security scan and the code comments in the file to determine whether Copilot generates the code snippet with the security problem (this process is explained further in Section 3.3.2). As a result, we obtained 452 code files with different contents, 272 from the Repository label and 180 from the Code label. Table 2 gives the type and number of code files. A curated dataset, comprising all the data collected during the research process, has been made available (Fu et al., 2024).
@@
(a) Code from the Repository label
(b) Code from the Code label
Figure 6. Application domains of the Copilot-generated code in our dataset
@@
L1 Python 143 134 277
L2 JavaScript 129 46 175
Total 272 180 452
3.3. Data Processing and Analysis
3.3.1. Data Processing
Step 4. CodeQL is a scalable static security analysis tool that is widely used in practice and allows users to analyze code and detect relevant weaknesses using predefined queries and test suites and supports for multiple languages (including Java, JavaScript, C++, C#, Go, and Python (GitHub, 2023a)). Before using CodeQL to scan the identified code snippets for security weaknesses, we needed to create a CodeQL database for the source code. Source code can be directly analyzed for interpreted languages like Python and JavaScript without being compiled into intermediate code. To speed up the scan, we stored 20 files in each database using the command line to generate the database needed for CodeqQL queries. It should be noted that generating a database for an exceptionally large number of files would increase the database compilation and scanning time, which is much longer than partitioning them into small databases. In total, we obtained 25 code databases available for CodeQL scanning.

### v3

@@ 2.4. Security Static Analysis
Detecting software vulnerabilities is critical to improve security and ensure quality  (Iannone et al., 2022). Vulnerabilities can be detected through static, dynamic, or hybrid analysis. Static analysis offers greater coverage and allows users to analyze programs without the need to execute them, but it can be unsound and imprecise  (Sui et al., 2020). On the other hand, dynamic analysis is more sound and precise (as it captures actual program behaviour at runtime) but can be incomplete and lacks coverage (Ernst, 2003). Due to their ease of use and low cost compared to dynamic analysis, static analysis tools are widely used for security analysis (Nunes et al., 2018; Huth and Nielson, 2019).
There are several security static analysis tools available for different programming languages and application types. OWASP and Snyk (OWASP, 2024; Snyk, 2024) provide a list of commonly used static analysis tools for security analysis. This includes tools like CodeQL, a general-purpose automatic analysis tool; FindBugs/SpotBugs for Java; ESLint for JavaScript programs; Bandit for Python; and GoSec for Go programs. Those tools have been widely used in previous security analysis research  (Siddiq et al., 2022; Tómasdóttir et al., 2018; Pearce et al., 2022).
Kaur et al. (Kaur and Nayyar, 2020) compared static analysis tools for vulnerability detection in analyzing C/C++ and Java source code. Tomasdottir et al. (Tómasdóttir et al., 2018) conducted an empirical study on ESLint, the most commonly used JavaScript static analysis tool among developers. Pearce et al. (Pearce et al., 2022) used CodeQL to scan security weaknesses in the generated Python and C++ code. Siddiq et al. (Siddiq and Santos, 2022) used Bandit to check Python code generated using a test dataset. Lisa et al. (Do et al., 2020) reported on users’ goals, motivations, and strategies when using static analysis tools.
These static analysis tools support different analysis algorithms and techniques. By using multiple tools for analysis, potential weaknesses in the code can be discovered from different perspectives and levels, avoiding omissions and improving the accuracy of the analysis. Our study first used CodeQL to scan the collected code snippets. CodeQL is an open-source tool that supports multiple languages, including Java, JavaScript, C++, C#, and Python. It can find weaknesses in a codebase based on known weaknesses/rules. In addition, to obtain more comprehensive analysis results, we supplemented the scan of code in different languages with static analysis tools (i.e., Cppcheck and Bandi) tailored to specific languages.
@@ 3. Research Design
In this section, we describe our research design in detail. In Section 3.1, we first define our Research Questions (RQs), followed by the process of collecting and filtering the code snippets generated by code generation tools such as Copilot in Section 3.2. We then explain the security analysis performed on the identified snippets and the process of filtering the raw results generated by static analysis tools in Section 3.3. Finally, in Section 3.4, we introduce the process of using Copilot Chat to fix the security weaknesses identified by the static analysis tools.
@@ 3.1. Research Goal and Questions
This study aims to understand the potential security weaknesses in AI-generated code produced in open source GitHub projects. We first collect code snippets generated by code generation tools such as Copilot from GitHub projects, which is our data source. It should be noted that it is not possible to access all the AI-generated code in GitHub projects, as there is no direct way to identify if part of a file was generated by code generation tools such as Copilot (i.e., source files do not contain any signatures to indicate if Copilot generates the code). However, we can identify many code snippets by searching the repository description and the comments provided in the code (see the details in Section 3.2.2).
We first used keyword-based search to collect projects and files containing code generated by Copilot and other AI code generation tools (CodeWhisperer and Codeium) from GitHub and filtered them manually. We subsequently analyzed the functionality and application domains of the collected code snippets to get a comprehensive understanding of the dataset. Next, we performed a static security analysis on the identified code.
After running the analysis and obtaining the analysis results, we manually checked the results to remove false positives reported by the static analysis tools. We then used CWEs to classify the filtered results for further analysis to answer our Research Questions (RQs). In addition, we used Copilot Chat (GitHub, 2024a) to fix code snippets containing security weaknesses. Copilot Chat is an interactive tool launched by OpenAI based on the GPT-4 model, designed to enable natural language-driven coding as part of Copilot. Developers can use Copilot Chat to perform tasks such as code analysis and fixing bugs. We used the warnings from static analysis tools as prompts and fed them to Copilot Chat in order to get the fixes. We then performed a security analysis on the fixed code to evaluate the effectiveness of those fixes.
To conduct our empirical study, we followed the guidelines of Easterbrook et al. (Easterbrook et al., 2008). The RQs, their rationale, and the research process of this study (see Figure 3) are detailed in the following subsections.
@@
Rationale: Copilot may produce code suggestions that developers accept, but these suggestions may include security weaknesses that could make the program vulnerable. The answer to RQ1 helps understand the frequency of security weaknesses developers encounter when using Copilot in production. We also include the code generated by other two popular AI code generation tools, CodeWhisperer and Codeium, in order to increase the generalizability of this research.
RQ2. What security weaknesses are present in the code generated by Copilot across different application domains?
Rationale: Copilot-generated code may contain security weaknesses (Pearce et al., 2022; Majdinasab et al., 2024), and developers should conduct a rigorous security review before accepting the generated code suggestions. Copilot’s documentation notes the following: “users of Copilot are responsible for ensuring the security and quality of their code” (GitHub, 2023). By identifying common or recurring weaknesses, developers can be more prepared to prevent, mitigate, or fix these security weaknesses. The answer to RQ2 can help developers better understand possible security weaknesses in the code generated by Copilot and other AI code generation tools, as well as the characteristics of code that contains certain types of security weaknesses, allowing them to be vigilant about the generated code before it is integrated into their code base.
RQ3. Can Copilot Chat help fix the security weaknesses found in the code generated by Copilot?
Rationale: The purpose of Copilot Chat is to assist developers by answering questions about their code or providing fixes for issues identified in the code. Copilot Chat is integrated with Copilot to form a powerful AI assistant that can help developers build at their speed of thought in the natural language of their choice (GitHub, 2023). It is claimed that Copilot Chat can assist in fixing security weaknesses. In this RQ, we utilize Copilot Chat to examine whether it can help fix the security weaknesses identified in the Copilot-generated code because most of the security weaknesses in RQ1 and RQ2 come from Copilot. Our approach is to utilize Copilot Chat in the loop by collect generated code →analyze the code for security weaknesses →check for fixes in Copilot Chat →re-analyze the code to check if the fix is correct. Answering RQ3 can help to understand the capability of Copilot Chat in fixing Copilot-generated code that contains security issues.
@@ 3.2. Data Collection and Filtering
We chose GitHub as the primary data source for answering our RQs. As the world’s largest code hosting platform, GitHub contains millions of public code repositories and offers access to a large number of code resources, allowing us to cover multiple programming languages and project types in our study (Cosentino et al., 2017). At the same time, thousands of developers in the GitHub community have shared their experiences of using Copilot in open-source projects. We identified code snippets generated by AI code generation tools from GitHub projects.
We chose Copilot as our main research subject as it is a well-known commercial instance of an AI pair programming tool that has gained increased popularity among developers since its launch in 2021. Copilot is one of the most popular general-purpose AI code generation tools and only second to ChatGPT in terms of overall AI assistance tools developers use in practice (Overflow, 2024). The security drawbacks of code generated by Copilot have been investigated in previous studies (Pearce et al., 2022; McDaniel, 2021; Dow, 2021). To increase the generalizability of the research, we also attempted to collect data from two popular AI code generation tools, CodeWhisperer and Codeium - both are general-purpose, AI-driven code generators. We did not collect data from specialized AI code generation tools, such as AlphaCode as it focuses on solving complex coding tasks rather than everyday code generation.
@@ 3.2.1. Code Snippets Collection
Step 1. We utilized GitHub REST API to automatically collect projects and files containing AI-generated code. We started with a pilot search on GitHub to formulate our search keywords. First, we used the names of the code generation tools, such as “GitHub Copilot” and “CodeWhisperer” as the basic search terms. It should be noted that many code snippets in the search results contained the “GitHub Copilot” keyword, but the keyword was not used to declare that the code snippets were generated by Copilot. Developers may use them to describe their experience of using Copilot to generate code or showcase information related to Copilot. These code snippets are not what we seek as they do not directly relate to the code generated by Copilot. By manually analyzing the search results, we found that restricting the search terms provided more accurate results. We decided to use keyword combinations such as {by, use, with} + {GitHub Copilot, CodeWhisperer, Codeium} as search keywords. This approach enabled us to focus more on the code generated using AI code generation tools (e.g., Copilot) rather than code snippets that contain other content related to the generation tools.
In addition, we further limited the types of code snippets during the search to Python and JavaScript. Both are dynamically typed languages where a full analysis can be conducted without the need for compilation. On the other hand, statically typed languages such as Java and C++ require the code to be compiled first to generate the database required for analysis. Our data collection method makes it difficult to properly compile individual code snippets, given that some code is unavailable when conducting the analysis.
We selected Python and JavaScript as they are currently the two most popular programming languages used by developers (GitHub, 2024b), and the most frequently used with GitHub Copilot (Zhang et al., 2023). The popularity of the two languages ensures that we can collect a substantial amount of generated code from GitHub projects, providing a significantly larger sample size and statistically meaningful results compared to other languages. We collected the search results under the filter Code with file names, file paths, and forks and stars of the repositories where the files were located. Considering that some projects declared using GitHub Copilot generated code in their README files or project descriptions provided in GitHub, we decided to retain the results from the filter Repo with project name, project path, project description, and forks and stars of the projects.
@@
Table 1 shows a breakdown of the search terms we used and the number of search results per language. The same search result may contain multiple keywords, meaning there are duplicate projects in the collected data. After removing duplications, we ended up with 3,589 search results, of which 3,141 were from the Code label and 445 from the Repository label.
Table 1. Search results based on different terms
Search Term Code (Py) Code (Js) Repo (Py) Repo (Js) Total
by GitHub Copilot 1320 1256 48 43 2667
use GitHub Copilot 1344 1384 84 93 2905
with GitHub Copilot 1532 1876 95 177 3680
by CodeWhisperer 39 44 0 1 84
use CodeWhisperer 41 46 4 2 93
with CodeWhisperer 68 39 4 0 111
by Codeium 85 73 0 1 159
use Codeium 31 26 0 1 58
with Codeium 99 69 0 2 170
Total 4559 4813 235 320 9927
@@ 3.2.2. Filtering Code Snippets
Step 2. After obtaining the data from the keyword searches, we further filtered them by not only considering the accuracy of the keywords but also by investigating the project’s documentation, code comments, and other metadata in the search results to determine whether they were generated by GitHub Copilot and other code generation tools. Additionally, since we aimed to obtain code snippets used in open-source GitHub projects, we manually excluded search results related to solving simple programming practice problems on platforms that provide foundational exercises for enhancing algorithmic thinking and coding skills, such as bisection lookup and fast sorting tasks from LeetCode. We consider these programming exercises to be more closely related to correctness than security.
@@
We begin by explaining the terminology used in data filtering: the search results under the Repository label are the projects that contain code files, and the search results under the Code label are individual code files. Those code files contain code snippets generated by Copilot. In filtering the projects, we followed three criteria including two inclusion criteria and one exclusion criterion. Inclusion Criterion 1: For search results under the Repository label, we identified projects that are fully generated by Copilot, as declared in the project description or the associated README file(s). We retained code files for Python and JavaScript. Inclusion Criterion 2: For search results under the Code label, we retained code files with comments showing the code generated by Copilot. Exclusion Criterion 1: We excluded code files used to solve simple programming practice problems. Simple programming practice problems mainly refer to those standard coding exercises on platforms such as LeetCode, which are often named BOJ-number. We provide examples for the three criteria in Figure 4, Figure 5, and Figure 6. As shown in the example in Figure 4 for the Repository label, we kept all the Python files. In the next example in Figure 5, we kept the entire file where the Copilot-generated code snippet was located. In Figure 6, the code snippet was removed as it was determined that the code just solved a simple algorithmic problem.
@@
We manually filtered all the data, labeling whether they were generated by Copilot and other code generation tools based on the three criteria mentioned above. Two authors independently completed this manual filtration process for all the data within a period of two weeks. The first author compared the labeling results by the two coders and calculated the level of agreement between them using Cohen’s Kappa coefficient (Cohen, 1960). Cohen’s Kappa coefficient was 0.84, which was higher than 0.8, indicating a high level of agreement between the two coders and ensuring a good accuracy of the labeling results. Besides, for those results where the two authors disagreed on the labeling, they resolved the disagreements through a negotiated agreement approach (Campbell et al., 2013) during a meeting. After manual filtration, we retained 116 projects under the Repository label and 335 code files under the Code label.
For the code files retained under the Repository label, we consider the entire code file as code generated by Copilot and other code generation tools. In other words, we assume that Copilot and other code generation tools generate all the code in the file because it was stated in the README file that it was all generated by Copilot and other code generation tools. For code files retained under the Code label, we know that the files contain code snippets, perhaps even just a few lines of code, generated by Copilot and other code generation tools. Instead of identifying the specific AI-generated code in this step, we combine the warning messages from the security analysis and the code comments in the file to determine whether Copilot and other code generation tools generate the code snippet with the security problem (this process is explained further in Section 3.3.2). As a result, our dataset consisted of 733 distinct code files, with 672 from Copilot, 38 from CodeWhisperer, and 23 from Codeium. Table 2 gives the type and number of code files. A curated dataset, comprising all the data collected during the research process, has been made available (Fu et al., 2024).
@@
L1 Python 190 229 419
L2 JavaScript 208 106 314
Total 398 335 733
Step 3. To understand the features of our dataset, we first conducted a statistical analysis of the stars and forks of the projects where AI-generated code was present. The distribution of stars and forks are shown in Figure 7 and Figure 8, respectively. We found that the projects containing AI-generated code often tend to be small and have low popularity in the GitHub community, possibly often developed by individuals or small teams. This finding is consistent with the recent study by Yu et al. about the characteristics of LLM-generated code on GitHub (Yu et al., 2024). We speculated that the low popularity may be correlated with the age of the project. Most of these projects are relatively new as Copilot itself is also a recent development. Despite the low popularity and user engagement of most projects, there are a few exceptions in the dataset, such as the project OnmyojiAutoScript, which received 1,672 stars and 505 forks.
In addition, we also counted the lines of code (LoC) of the collected code files. Figure 9 shows the distribution of LoC in different languages (Python and JavaScript respectively) and then the LoC of all code files (Python + JavaScript) in the dataset. We found that the median LoC of code is 74, while the average reaches 181. Although the files obtained from GitHub are relatively small in size, they are still significantly larger than the size of code generated using designed scenario prompts. For example, when Majdinasab et al. (Majdinasab et al., 2024) replicated the experiment by Pearce et al. (Pearce et al., 2022), the median LoC of code generated by their designed prompts through Copilot was 42, with an average of 40. While using the SecurityEval dataset (Siddiq and Santos, 2022) as prompts to test Copilot, the LoC of generated code is concentrated between 5 and 26. This indicates that our collected dataset has an advantage in terms of the size of the generated code.
Figure 7. Distribution of stars of the projects where AI-generated code was present
Figure 8. Distribution of forks of the projects where AI-generated code was present
Figure 9. Distribution of LoC of the files where AI-generated code was present
We also analyzed the application domains of the code snippets in our dataset. To understand the application domains, we classified the code in the dataset based on the project description and the specific functions of the code files. Two authors independently classified all code snippets within a period of ten days. If their categorization results diverged, the two authors discussed until they reached a consensus. In total, we categorized the generated code into six application domain categories: Game, Web Application, Utility Tool, AI Application, Network Communication, and Others. Table 3 shows the descriptions of application domains and the categorization of code snippets in different programming languages.
Table 3. Application domains of the files where AI-generated code was present
Category Description Python JavaScript
Utility Tool Involves scripts, command-line tools, and other programs that simplify daily tasks. 39.0% 30.1%
Web Application Code related to front-end, back-end development, and server-side operations. 15.6% 49.8%
AI Application Code for machine learning, deep learning, natural language processing, or other intelligent applications. 15.6% 2.8%
Game Contains code related to game development, using engines such as Unity, Unreal, and Godot. 15.9% 8.5%
Network Communication Involves code for implementing network protocols, client/server communication. 8.2% 5.7%
Others Code that does not fit into the above categories. 5.4% 2.8%
3.3. Data Analysis and Results Filtering